hribf_logo
home printable Archive hribf Physics Contact
   


RA2. Accelerator System Status

ORIC Operations and Development (B. A. Tatum)

ORIC operated for much of the reporting period, providing proton, deuteron, and alpha beams for the production of 664 hours of RIB on target for the experimental program. The only major problem encountered during the period that required significant downtime was a small water leak on the coaxial magnetic extraction channel. This leak was temporarily repaired by an internal sealant, but ultimately the channel had to be removed from the machine for replacement of the insert coil. Removal and repair of any extraction channel requires 3-6 weeks of downtime due to the complexity of the work and radiological conditions.

New harmonic coil power supplies have been purchased to replace aging supplies that have been a chronic source of instabilities in recent months. ORIC contains four sets of harmonic coils, but only the #1 and #4 (or valley) coils are required for beams that we presently accelerate. The new #4 power supplies are now in service and the #1 power supplies will be received and installed in the coming quarter. To help maintain good vacuum in ORIC, a new 20" cryopump has been procured and will be installed in the near future. This pump replaces a failed unit located on the south end of the resonator which provides supplemental pumping to the three diffusion pumps. We have also ordered a new 4648 rf power amplifier tube from Burle Industries to ensure many more years of operation.

ORNL is in the process of replacing our almost 50-year-old substation circuit breakers. The modern breakers, which are expected to be received and installed prior to the end of the fiscal year, should improve facility reliability and personnel safety.

RIB Injector Operations and Development (P.E. Mueller)

During the period from 1 January 2008 to 30 June 2008, the 25 MV Tandem Electrostatic Accelerator delivered beams of

  • 13 Mpps [5.31 MV 3+ terminal gas stripped] 21.2-MeV 17F,
  • 8 Mpps [3.54 MV 2+ terminal gas stripped] 10.8-MeV 17F, and
  • 13 Mpps [4.73 MV 2+ terminal gas stripped] 14.3-MeV 17F to the H2 gas target in front of the Daresbury Recoil Separator and
  • 40 Mpps [21.35 MV 7+ / 9+ terminal foil / post foil stripped] 171-MeV pure 17F to Beam Line 21.

    The 17F beams were produced via the 16O(d, n)17F reaction by bombarding a fibrous HfO2 target coupled to a Kinetic Ejection Negative Ion Source (KENIS) with 3-5 uA of 44-MeV 2H from the Oak Ridge Isochronous Cyclotron (ORIC).

  • Tandem Operations and Development (M. Meigs)

    The Tandem Accelerator was operated for more than 2700 hours since the last report. The machine ran at terminal potentials of 2.02 to 22.87 MV and the stable beams 1H, 12C, 16,17,18O, 19F, 31P, 35Cl, 56Fe, 58Ni, 107Ag, and 118,120Sn were provided. Radioactive beams of 17,18F accounted for more than 660 hours of beam on target. Conditioning was done for about 24 hours to return the machine to operation at 23.5 MV after having the tank open for maintenance. Two tank openings were required during this period; the first for scheduled maintenance and the second to repair a shorting rod contact that had been incorrectly installed in the first tank opening.



    Home printable Archive hribf Physics ornl contact